OpenStudy (anonymous):

can (x^2 -2)^2 be differentiated?

OpenStudy (anonymous):

directly i mean, without expanding

OpenStudy (anonymous):

Yes, (x^2-2)^2= x^4-4x^2+4 When differentiated this will be: 4x^3-8x

OpenStudy (anonymous):

i remenber in some cases only differentiating linear functions is possible,but cant recall in what cases you cant differentiate x something with a square

OpenStudy (raden):

use chain's rule y=(x^2-2)^2 let u=x^2-2 du/dx=.... y=u^2 dy/du=.... dy/dx=du/dx * dy/du